# Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition

### v1

@@ 4.1 Experimental Setup
To ensure a fair comparison, all experimental setups are consistent across all methods. Specifically, for training-free methods, we randomly select 512 samples from the WikiText-2 dataset as calibration data. For training-based methods, all models are fine-tuned on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32. We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$. All our experiments are conducted on a single NVIDIA A100 GPU with mixed precision enabled. Further experimental details can be found in the Appendix A.3.
Model Method Ratio WikiText-2 PTB C4 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B Dense 0% 5.25 23.73 8.36 32.6 80.85 73.88 61.22 50.43 80.58 35.88 59.35
LaCo 21.1% 244.53 524.26 243.98 20.2 35.17 57.54 26.23 24.66 52.77 24.29 34.41
SliceGPT 20% 9.04 107.65 38.17 19.2 51.05 59.43 34.84 25.26 60.72 23.25 39.11
ASVD 20% 19.28 171.54 27.93 19.8 55.68 55.96 35.84 24.49 66.21 24.19 40.31
Taco-SVD 20.3% 26.42 77.98 28.88 21.4 51.47 68.19 42.82 35.24 66.59 25.83 44.51
LLaMA2-7B Dense 0% 5.47 20.83 7.29 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LaCo 18.1% 39.43 93.98 71.77 26.2 48.36 59.19 41.93 32.42 68.72 24.02 42.98
SliceGPT 21.5% 8.63 99.25 40.7 22 54.34 60.62 37.4 27.99 63.38 23.18 41.27
ASVD 20% 9.7 88.39 12.46 27.8 58.41 61.09 46.97 32.42 71.05 25.09 46.12
Taco-SVD 21.6% 16.12 44.09 23.13 24 53.59 62.9 42.94 32.76 71.06 22.88 44.30
LLaMA-2-13B Dense 0% 4.88 28.92 6.99 32.4 73.27 69.61 59.72 45.56 78.78 30.02 55.62
LaCo 19.5% 27.26 59.14 27.42 27.6 52.25 62.98 42.52 33.36 69.91 25.32 44.85
SliceGPT 21.5% 7.44 99.77 37.65 29.4 58.92 64.8 38.96 31.74 64.42 23.65 44.56
ASVD 20% 8.07 101.1 11.86 28.8 63.55 62.59 46.12 32.68 73.72 24.65 47.44
Taco-SVD 20.6% 11.1 56.57 14.76 26.4 60.62 66.22 46.55 35.07 73.11 24.22 47.46
LLaMA3.1-8B -Instruct Dense 0% 7.21 12.33 12.06 33.8 81.9 73.72 59.1 51.71 79.87 39.46 59.94
LaCo 19% 753.98 1577.44 807.12 26 48.78 64.56 33.2 29.61 65.07 30.25 42.50
SliceGPT 20% 52.78 460.73 324.71 14.6 43.48 51.38 29.6 22.78 58.38 21.78 34.57
ASVD 20% 2443.99 4469.54 2842.23 11.8 28.37 48.62 26.63 19.2 54.84 18.79 29.75
Taco-SVD 20.1% 37.86 63.97 42.25 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
Table 1: Zero-shot performance of Taco-SVD and training-free baselines under 20% compression ratio. Results are reported on three language modeling datasets (perplexity) and seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
Model Method Ratio WikiText-2 PTB C4 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B Dense 0% 5.25 23.73 8.36 32.6 80.85 73.88 61.22 50.43 80.58 35.88 59.35
LoRA 25% 11.56 34.05 14.86 24.6 68.73 70.72 51.18 40.1 71.93 26.37 50.52
LS-FFN 25.2% 16.12 59.98 18.4 21.8 66.88 68.03 49.92 37.8 70.35 27.2 48.85
LS-Layer 24.1% 14.67 54.31 16.61 21.6 68.77 70.09 45.02 39.25 72.25 26.83 49.12
Taco-SVD 24.3% 11.62 44.14 14.85 26.2 69.28 70.32 51.7 40.87 73.07 27.17 51.23
LLaMA-2-7B Dense 0% 5.47 20.83 7.29 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LoRA 25% 11.7 35.39 13.95 26.2 57.79 63.06 48.15 33.28 71.16 24.15 46.26
LS-FFN 25.7% 13.94 37.43 14.61 24.6 55.98 62.51 48.32 33.28 71 25.19 45.84
LS-Layer 24% 11.57 32.62 12.91 25.4 59.85 61.88 50.21 35.15 72.91 23.85 47.04
Taco-SVD 24.2% 9.59 30.38 12.22 28.4 60.52 64.56 50.32 36.01 71.76 24.85 48.06
LLaMA-2-13B Dense 0% 4.88 28.92 6.99 32.4 73.27 69.61 59.72 45.56 78.78 30.02 55.62
LoRA 25% 8.46 58.15 11.35 26.6 63.05 67.4 53.47 38.4 73.5 25.33 49.68
LS-FFN 25.4% 8.11 76.3 11.22 28.4 61.07 67.17 54.04 38.23 73.29 25.49 49.67
LS-Layer 24.4% 8.23 77.04 11.09 29 62.42 67.25 54.25 39.76 74.21 26.06 50.42
Taco-SVD 24.2% 7.52 38.31 10.42 28.6 63.3 67.96 54.63 38.74 73.88 27.27 50.63
LLaMA3.1-8B -Instruct Dense 0% 7.21 12.33 12.06 33.8 81.9 73.72 59.1 51.71 79.87 39.46 59.94
LoRA 22.3% 14.56 21.24 21.8 27.6 69.65 69.38 50.54 41.3 72.69 30.18 51.62
LS-FFN 22% 16.79 24.76 24.14 26.4 70.24 70.17 48.97 40.53 72.03 30.25 51.23
LS-Layer 21.7% 14.27 20.85 20.79 27.2 73.53 69.38 51.22 42.66 73.39 31.56 52.71
Taco-SVD 21% 14.13 20.99 21.73 28.2 72.01 69.93 51.5 43.26 73.5 31.26 52.81
@@
Table 2: Performance comparison of Taco-SVD + FT and training-based baselines under 25% compression ratio. LS-FFN and LS-Layer denote two variants of LLM-Streamline. Results are reported on three language modeling datasets (perplexity) and seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
4.2 Results
We conducted comprehensive experiments to evaluate Taco-SVD from four key perspectives: (a) zero-shot compression performance on language generation and reasoning tasks, (b) effectiveness as an initialization method compared to training-based approaches, (c) robustness at high compression ratios, and (d) compression efficiency and inference speed on real hardware.
Training-Free Compression
To evaluate the generalizability of Taco-SVD, we compared its performance with training-free baselines under approximately 20% compression ratio across four models with varying scales and architectures: Mistral-7B, LLaMA2-7B, LLaMA2-13B, and LLaMA3.1-8B-Instruct. The results, summarized in Table 1, demonstrate that Taco-SVD consistently achieves strong performance across all models and tasks, maintaining competitive perplexity and accuracy even on challenging architectures like LLaMA3. Notably, Taco-SVD exhibits significantly improved robustness compared to baselines, effectively addressing the variability in sensitivity to pruning across diverse architectures.
Training-based Compression
In this section, we further demonstrate that Taco-SVD can provide better initialization for fine-tuning, enabling faster and more accurate recovery of the compressed model compared to other methods. To ensure that the number of trainable parameters remains approximately the same across all methods, we retain only 10% of the parameters in each redundant layer and set only the parameters within these redundant layers to be trainable. As shown in Table 2, Taco-SVD consistently outperforms LoRA and LLM-Streamline across different LLMs. In addition, we plot the validation loss curves during the training process for each method in the Appendix A.4, we can observe that Taco-SVD has better initialization, leading to more stable and faster convergence during fine-tuning.
Performance at High Compression Ratio
To evaluate the performance and stability of our proposed method at high compression ratios, we compare Taco-SVD with the baselines on LLaMA2-7B at compression ratios of 25%, 30%, 35% and 40%. The perplexity (PPL) on WikiText-2, PTB and the average accuracy across all reasoning benchmarks are shown in Figure 3, with detailed results for each benchmark presented in the Appendix A.5. We make the following three observations: (a) Most layer removal methods cannot tolerate high compression ratios, with performance dropping sharply at higher compression levels. (b) Compared to training-free baselines, Taco-SVD consistently achieves better accuracy across all compression ratios. (c) Fine-tuning Taco-SVD significantly recovers the accuracy of compressed models, even at high compression ratios.
Figure 3: Zero-shot performance of LLaMA2-7B compressed by Taco-SVD and baselines under 20% to 40% compression ratio. Top: Perplexity on WikiText-2 (left) and PTB (right). Bottom: Average accuracy on seven reasoning benchmarks.
Compression Costs and Acceleration Effect
Taco-SVD can not only compress LLMs at very low cost but also improve inference efficiency on real hardware. To demonstrate this, we measure the compression time and peak memory consumption of Taco-SVD when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio. These metrics, summarized in Table 3, indicate that Taco-SVD achieves a good balance between compression time and memory consumption. Additionally, to assess the acceleration effect of Taco-SVD, we compared the throughput (tokens generated per second) of original LLaMA2-7B and its compressed version by layer removal and Taco-SVD with different sequence lengths and batch sizes. As shown in Figure 4, Taco-SVD consistently ensures an enhancement in the generation speed and achieves an acceleration effect comparable to pure layer removal. It is important to note that although Taco-SVD retains a small portion of parameters in the redundant layers to mitigate the internal inconsistencies introduced by layer removal, this structure is extremely low-rank, resulting in negligible inference latency while maintaining task performance.
@@ 4.2 Results
Compression Time(h)
Memory
Consumption(MiB)
LaCO 0.05 45046
ASVD 5.5 17632
SliceGPT 0.6 18195
Taco-SVD 0.16 29050
Table 3: Compression time and peak memory usage of Taco-SVD and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
@@
Figure 4: Throughput of LLaMA2-7B and Taco-SVD compressed model under 25% compression ratio on a single A100 GPU. Top: Throughput across different sequence lengths (batch size = 32). Bottom: Throughput across different batch sizes (sequence length = 32).
@@ 4.3 Ablation Study
We conducted ablation studies to investigate the impact of calibration dataset and the number of calibration data used for singular value gradient attribution. Table 4 summarizes the results of Taco-SVD when compressing LLaMA3.1-8B-Instruct under 20% compression ratio using different calibration datasets and varying numbers of calibration data from WikiText-2. The results demonstrate that Taco-SVD is robust to both the choice of calibration dataset and the quantity of calibration data. Notably, even with as few as 64 calibration samples, Taco-SVD achieves competitive performance, highlighting its efficiency in leveraging minimal data for effective compression. We report the detailed results in the Appendix A.6.
@@
Table 4: Comparison of Taco-SVD using different types and amounts of calibration data for compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
@@
As detailed in Section 3.4, Taco-SVD employs a layer-wise backward compression strategy to compress redundant layers. This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers. Table 5 shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
Pruning Strategy WikiText-2 PTB
Average
Accuracy
Compression
Time(h)
One-shot Pruning 37.86 63.97 47.12 0.16
Iterative Pruning 38.39 72.18 47.13 0.22

### v2

@@ 4.1 Experimental Setup
For all baselines, we randomly sample 512 instances from the WikiText-2 dataset as the representation evaluation dataset. For pruning-based with post-training compensation baselines, all models are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32. Further experimental details can be found in Appendix A.3.
Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
Method
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
Original 5.25 23.73 0.59 5.47 20.83 0.53 4.88 28.92 0.56 7.21 12.33 0.60
ShortGPT 23.95 80.70 0.46 18.45 57.58 0.43 27.26 59.14 0.44 140.47 262.47 0.44
LaCo 244.53 524.26 0.34 39.43 93.98 0.43 38.27 125.98 0.42 753.98 1577.44 0.43
SliceGPT 9.04 107.65 0.39 8.63 99.25 0.41 7.44 99.77 0.45 52.78 460.73 0.35
Ours 18.42 77.98 0.46 16.12 44.09 0.44 11.1 56.57 0.47 37.86 63.97 0.47
Table 1: Perplexity($\downarrow$) of GRASP and pruning-based without post-training compensation baselines on the WikiText-2 and PTB datasets, as well as the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio. The best performance is marked in bold.
Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
Method
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
PPL $\downarrow$
Acc $\uparrow$
Average
Original 5.25 23.73 0.59 5.47 20.83 0.53 4.88 28.92 0.56 7.21 12.33 0.60
LoRA 11.56 34.05 0.51 11.7 35.39 0.46 8.46 58.15 0.50 14.56 21.24 0.52
LS-FFN 16.12 59.98 0.49 13.94 37.43 0.46 8.11 76.3 0.50 16.79 24.76 0.51
LS-Layer 14.67 54.31 0.49 11.57 32.62 0.47 8.23 77.04 0.50 14.27 20.85 0.53
Ours* 11.62 44.14 0.51 9.59 30.38 0.48 7.52 38.31 0.51 14.13 20.99 0.53
@@
Table 2: Perplexity($\downarrow$) of GRASP (Ours* denotes the version with post-training compensation) and pruning-based with post-training compensation baselines on the WikiText-2 and PTB datasets, as well as the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 25% compression ratio. LS-FFN and LS-Layer denote two variants of LLM-Streamline. The best performance is marked in bold.
Ratio Method WikiText-2 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA GSM8K Average Percentage
0% Original 5.68 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.09 0.46 100.0
20% FWSVD 1727 0.15 0.31 0.50 0.26 0.23 0.56 0.21 0.00 0.28 60.8
ASVD 11.14 0.25 0.53 0.64 0.41 0.27 0.68 0.24 0.04 0.38 82.6
SVD-LLM 7.94 0.22 0.58 0.63 0.43 0.29 0.69 0.24 0.05 0.39 84.7
Ours 14.79 0.22 0.52 0.64 0.43 0.32 0.70 0.24 0.03 0.39 84.7
Ours* 10.19 0.24 0.59 0.63 0.5 0.35 0.73 0.25 0.04 0.42 91.3
30% FWSVD 20127 0.17 0.26 0.49 0.26 0.22 0.51 0.19 0.00 0.26 56.5
ASVD 51 0.18 0.43 0.53 0.37 0.25 0.65 0.21 0.00 0.33 71.7
SVD-LLM 9.56 0.20 0.48 0.59 0.37 0.26 0.65 0.22 0.03 0.35 76.1
Ours 30.48 0.19 0.42 0.62 0.39 0.28 0.64 0.23 0.02 0.35 76.1
Ours* 12.63 0.24 0.54 0.64 0.46 0.32 0.69 0.24 0.04 0.40 87.0
40% FWSVD 18156 0.16 0.26 0.51 0.26 0.22 0.53 0.21 0.00 0.27 58.7
ASVD 15358 0.13 0.28 0.48 0.26 0.22 0.55 0.19 0.00 0.26 56.5
SVD-LLM 13.11 0.19 0.42 0.58 0.33 0.25 0.60 0.21 0.02 0.33 71.7
Ours 62.34 0.18 0.37 0.57 0.35 0.27 0.61 0.21 0.01 0.32 69.6
Ours* 16.25 0.22 0.49 0.63 0.43 0.3 0.68 0.23 0.02 0.38 82.6
50% FWSVD 24391 0.12 0.26 0.50 0.26 0.23 0.53 0.20 0.00 0.26 56.5
ASVD 15358 0.12 0.26 0.51 0.26 0.22 0.52 0.19 0.00 0.26 56.5
SVD-LLM 23.97 0.16 0.33 0.54 0.29 0.23 0.56 0.21 0.00 0.29 63.0
Ours 102.58 0.13 0.29 0.53 0.28 0.23 0.53 0.20 0.01 0.28 60.9
Ours* 23.69 0.18 0.4 0.56 0.35 0.26 0.61 0.21 0.02 0.32 69.6
[…]
Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
Method
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
PPL $\downarrow$
(WikiText-2)
Acc $\uparrow$
Average
Original 5.25 0.59 5.68 0.52 5.47 0.53 7.21 0.60
FWSVD 6357 0.32 1727 0.32 2360 0.31 3256.7 0.29
ASVD 19.28 0.4 11.14 0.44 9.70 0.46 2443.99 0.30
SVD-LLM 10.21 0.41 7.94 0.44 8.50 0.47 - -
Ours 18.42 0.45 14.79 0.44 16.12 0.44 37.86 0.47
Ours* 11.62 0.51 10.19 0.47 9.59 0.48 14.13 0.53
Table 4: Perplexity($\downarrow$) of GRASP and SVD-based baselines on the WikiText-2 datasets and the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio. "†" indicates that we refer to the results in the original paper. The best performance is marked in bold.
@@
Single-Doc
QA
Multi-Doc
QA
Summarization
Task
Few-shot
Learning
Synthetic
Task
Code
Completion
@@
LLaMA3.1-8B-Instruct 12.26 16.16 22.19 42.52 9.49 53.91 26.09
GRASP* 21.69 24.41 20.03 45.84 9.67 35.92 26.26
Table 5: Evaluations of LLaMA3.1-8B-Instruct and its compressed version of GRASP* under 20% compression ratio on LongBench benchmarks. The average results across six task categories are shown in this table, with additional details provided in the Appendix A.7.
4.2 Comparison with Pruning-based LLM Compression Methods
Post-training compensation plays a crucial role in restoring the performance of compressed LLMs. To ensure a fair comparison, we compare GRASP with two categories of pruning-based compression methods. (1) Pruning without post-training compensation, including two layer removal methods, ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024), and state-of-the-art structured pruning method: SliceGPT (Ashkboos et al., 2024). (2) Pruning with post-training compensation: ShortGPT with LoRA (Hu et al., 2022) and LLM-Streamline (Chen et al., 2024). We provide a detailed comparison of these pruning-based LLM compression methods in Appendix A.4.
Pruning without post-training compensation
We present the perplexity on WikiText-2 and PTB datasets, as well as the average accuracy on common sense reasoning benchmarks for different methods in Table 1. The results demonstrate that GRASP consistently achieves strong performance across all models and tasks. Specifically, in common sense reasoning tasks, GRASP surpasses SliceGPT by 15% in accuracy on Mistral-7B and by 34% on LLaMA3.1-8B-Instruct. For language generation tasks, while SliceGPT achieves lower perplexity on the WikiText-2 dataset, GRASP consistently outperforms the baselines on the PTB dataset. We hypothesize this discrepancy arises from the potential overfitting problem of baselines to the calibration dataset. Notably, GRASP exhibits significantly improved robustness compared to baselines, effectively addressing the variability in sensitivity to pruning across diverse models. Detailed results for each common sense reasoning benchmark are provided in Appendix A.5.
Pruning with Post-Training Compensation
We further demonstrate that GRASP provides a better initialization for post-training compensation, enabling more accurate recovery of the compressed model compared to other methods. To ensure a fair comparison, we constrain the number of trainable parameters to remain approximately the same across all methods by retaining only 10% of the parameters in each redundant layer and allowing only these parameters to be trainable. As shown in Table 2, GRASP consistently outperforms LLM-Streamline across different LLMs. Furthermore, we present the validation loss curves for each method during post-training in Appendix A.6, which show that GRASP GRASP enables more stable and faster convergence. More Detailed results are provided in Appendix A.5.
4.3 Comparison with SVD-based LLM Compression Methods
In this section, we compare the performance of GRASP with state-of-the-art SVD-based LLM compression methods—FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024)—from two distinct perspectives: (1) performance under different compression ratios and (2) performance across different LLMs.
Performance under Different Compression Ratios.
To evaluate the performance and stability of our proposed method under various compression ratios, we compare GRASP with the SVD-based baselines on LLaMA-7B under compression ratios ranging from 20% to 50%. Table 3 summarizes the results for different methods. The results demonstrate that our proposed GRASP consistently outperforms the baseline methods on most benchmarks, except for a slightly higher perplexity on WikiText-2 compared to SVD-LLM. Specifically, GRASP retains more than 91% of the original performance at a 20% compression ratio and 87% under the compression ratio of 30%. More importantly, with fast and resource-efficient post-training, GRASP enables rapid accuracy recovery, achieving 70% of the original model performance even at a 50% compression ratio.
Performance on different LLMs.
To evaluate the generalizability of GRASP across different LLMs, we compare its performance with the SVD-based baselines under 20% compression ratio on four models from two distinct LLM families-LLaMA 2-7B, LLaMA 2-13B, LLaMA 3.1-8B-Instruct and Mistral-7B. The results, presented in Table 4, indicate that while existing methods achieve competitive performance on the LLaMA 2 series, they exhibit a sharp performance drop on other architectures, particularly on LLaMA 3. In contrast, GRASP consistently outperforms baseline methods across all architectures and demonstrates greater stability across different LLMs.
4.4 Performance in Long-Form Text Processing
In this section, we evaluate the performance of GRASP on LongBench (Bai et al., 2023) to assess its performance on long-form reasoning and complex generative tasks. The average performance across six task categories: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion are summarized in Table 5, with detailed results provided in the Appendix A.7. As shown, the compressed version by GRASP under 20% compression ratio still maintains strong performance and even surpasses the original LLaMA3.1-8B-Instruct model in single-document QA, multi-document QA, and few-shot learning tasks. This demonstrates the effectiveness of GRASP in rapidly recovering accuracy and generative capability through post-training compensation after pruning.
4.5 Compression Costs and Inference Efficiency of GRASP
GRASP can compress LLMs at very low cost while improve the inference efficiency on real hardware. To evaluate its acceleration effect, we compared the throughput (tokens generated per second) of original LLaMA2-7B and its compressed version by GRASP with different sequence lengths and batch sizes. As shown in Figure 3, GRASP consistently consistently improves generation speed and achieves an acceleration effect comparable to pure layer removal. It is important to note that although GRASP retains a small portion of parameters in the redundant layers to mitigate the internal inconsistencies introduced by layer removal, this structure is extremely low-rank, resulting in negligible inference latency while maintaining task performance. Additionally, we also measure the compression time and peak memory consumption of GRASP when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio. The detailed results are provided in Appendix A.8.
@@ 4.5 Compression Costs and Inference Efficiency of GRASP
Figure 3: Throughput of LLaMA2-7B and GRASP compressed model under 25% compression ratio on a single A100 GPU. Top: Throughput across different sequence lengths (batch size = 32). Bottom: Throughput across different batch sizes (sequence length = 32).
@@ 4.6 Ablation Study
We conducted ablation studies to investigate the impact of representation evaluation dataset and the number of data used for singular value gradient attribution. Table 6 summarizes GRASP’s performance when compressing LLaMA3.1-8B-Instruct using different representation evaluation datasets and varying numbers of data from WikiText-2. The results demonstrate that GRASP is robust to the choice of both the representation evaluation dataset and the data number. Notably, GRASP achieves competitive performance with just 64 samples, highlighting its ability to perform effective compression with minimal data. Detailed results can be found in Appendix A.9.
@@
Table 6: Comparison of GRASP using different types and amounts of data for compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
@@
As detailed in Section 3.4, GRASP employs a layer-wise backward compression strategy to compress redundant layers. This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers. We present the ablation results in Table 12 in Appendix A.10, which shows that both approaches achieve similar performance, with one-shot pruning being more efficient.